It’s Five​ O’clock Somewhere, LLC manufactures beverage containers. The company is anxious to produce and sell a new beverage container designed to keep beverages cool for up to 2 hours. The container will sell for​ $3 each. Enough capacity exists in the​ company’s plant to produce​ 14,000 of the new containers each month.​ $0.40 from every sales dollar is contribution margin. Fixed costs associated with the container would total​ $12,000 per month in the existing facility.

The​ company’s Marketing Department predicts that if demand for the new container exceeds the​ 14,000 containers that the company is able to produce in its current facility that additional manufacturing space can be rented from another company at a fixed cost of​ $2,000 per month. The rented facility has a production capacity of​ 8,000 units per month. The variable cost percentage in the rented facility would equal​ 70% due to somewhat less efficient operations than in the​ company’s current facility.

Which of the following statements is​ correct?

A. If the company produces less than​ 14,000 containers in its current​ facility, the total fixed costs will be less than ​$12,000.

B. The maximum monthly operating income that the company could make with the two facilities is ​$19,600.

C. To make a monthly target profit of​ $5,000, a total of​ 16,445 containers must be produced and sold.

D. As long as the​ company’s monthly target profit is less than or equal to​ $5,200, it will not need to rent the additional facility.

E. The monthly breakeven point in the rented facility is greater than the monthly breakeven point in the existing facility.
Answer #1

In the existing case,

the capacity = 14000 units , selling price = $ 3 per unit and contribution margin is 0.4 per dollar of sales i.e for $ 3 of sales price per uniit, contribution margin per unit = $ 3* 0.4 = $ 1.2. and total fixed cost in existing case = $12000.

Thus contribution margin in existing case = $ 1.2.

Existing Full capacity, Contribution margin amount = Contribution margin per unit * total units = $ 1.2 * 14000  

Contribution margin = $ 16,800

Thus operating income in existing case = Contribution margin - Fixed cost = $ 16800 - 12000 = $ 4800.

Proposed case:

Aditional increase in capacity = 8000 units, Additional Fixed cost = $ 2000 , and contribution margin per unit = 30% of sales price ( as varaible cost is 70% which means the contribution margin % is 30%.

Contribution margin per unit in proposed case = Sales price per unit * Rate = $ 3 * 30% = $ 0.90 per unit.

Thus the existing case, contribution margin = Units 8 contribution margin per unit = 8000 * 0.90

Contribution margin amount = $ 7200 , Fixed cost = $ 2000

Thus Operating income = Contribution margin amount - Fixed cost = $ 7200 - $ 2000 = $ 5200.

Combined :

Thus at full capacity the operating income = Existing + Proposed = $ 4800 + $ 5200 = $ 10,000.

To make the monthly profit of $ 5000, the existing capacity will not be enough, as it can maximum generate $4800 profit in the case, so additional facility needs to be utilised which involves additional cost.

So for the remaining profiit of $ 200 ( $ 5000 - 4800 ) , the new facility needs to be taken,

Desired sales units = ( desired profit + fixed cost ) / contribution margin per unit

Desired sales units = ( 200 + 2000 ) / 0.9 = 2200 / 0.9 = 2444.44 = 2445 units.

So total sale units to get profit of $ 5000 = 14000 units + 2445 units = 16445 units.

Conclusion :

Option A is incorrect as the fixed cost does not changes with the change in units of sales, even if the company produce less than 14000 units it will still incurr the same fixed cost of $ 12000.

Option B is incorrect as the maximum operating profit it can make is $ 10000 as calculated above

Option D is incorrect as in existing case, it can generate maximum profit of $ 4800 if operated fully, its not $ 5200 but $ 4800.

Option E is incorrect as break even depends upon the fixed cost higher the fixed cost higher the break even point, fixed cost in existing case is much more than proposed rented facility, so rented facility break even will be lower than existing.

Only the correct option is Option -C i.e to make a monthly target profit of​ $5,000, a total of​ 16,445 containers must be produced and sold as calculated above.

Thus the correct option is ---------------C i,e To make a monthly target profit of​ $5,000, a total of​ 16,445 containers must be produced and sold.

Feel free to comment doubts if any..